OpenStudy (anonymous):

Prove the following identity: cos2θ = cos2θ − sin2θ = 2cos2θ − 1 = 1 − 2sin2θ

OpenStudy (cwrw238):

?? you sure you've got that right??

OpenStudy (solomonzelman):

Way too many equal signs .

OpenStudy (cwrw238):

i think some of those should contain an exponent - like sin^2θ

OpenStudy (solomonzelman):

yes, and the others are pluses

OpenStudy (solomonzelman):

I would think it is \(\normalsize\color{black}{ \cos2θ + \cos2θ − \sin2θ + 2\cos2θ − 1 = 1 − 2\sin2θ}\)

OpenStudy (solomonzelman):

Depends on which `=` is a `+`. (Where the user forgot to use `SHIFT` )

OpenStudy (solomonzelman):

@Katieholmes please re-write the question.

OpenStudy (midhun.madhu1987):

Ambiguity in question...

OpenStudy (solomonzelman):

@Katieholmes , are you there? Please re-write the question.

OpenStudy (anonymous):

No I rechecked it and that is the question. Ill add a screenshot

OpenStudy (cwrw238):

I can think of a way to do this if we assume the identity cos ( A + B) = cos A cos B - sinA sinB but i'm not sure if thats acceptable

OpenStudy (cwrw238):

you'll also need to apply sin^2A + cos^2 A = 1

OpenStudy (lochana):

cos2x = cos(x +x)= cosx.cosx - sinx.sinx = (cosx)^2 - (sinx)^2 therefor cos2x = (cosx)^2 - (sinx)^2 {x = theta} and cos2x = {1 - (sinx)^2) - (sinx)^2 = 1 - 2(sinx)^2 and cos2x = (cos2x)^2 - {1 - (cosx)^2} = 2(cosx)^2 - 1

OpenStudy (cwrw238):

- yes thats basically the way I suggested

OpenStudy (anonymous):

yea i know what you guys are talking about. It confused my because of all the different = signs and I though it just didnt seem right

OpenStudy (cwrw238):

yes also it was confusing because you didnt right the 'squared' bit as '^2'

OpenStudy (solomonzelman):

Katie, talk to your teacher get the correct question, and then ask it in a different question.